X + y = -1. -6x + 2y = 14
Answer by addingup(3677)

Multiply the 1st equation times -2 and you get:
-2x+-2y = 2 now add this from the other equation:
-6x+2y = 14
+
-2x-2y = 2
----------
-8x 0y = 16
-8x = 16
x = -2
Now that you have x you can solve for y:
x+y = -1
-2+y = -1
y = 1
------------------------------
Check - plug my answers into the original equations:
x+y = -1; -2+1 = -1 correct
-6x+2y = 14
-6(-2)+2(1) = 14
12+2 = 14 also correct